TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-18 22:46:00
subject: 3\03 Pt-1 ISS On Orbit Status 03-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\03 ISS On-Orbit Status 03-03-2003
Part 1 of 2

ISS On-Orbit Status 3/3/03

All ISS systems continue to function nominally, except as noted 
previously or below. Space Day 100 for Expedition 6 (since launch on 
STS-113 on 11/23/02). It will be 100 days next Thursday, 3/6, since 
Station command was transferred from Expedition 5's Valery Korzun to 
Kenneth Bowersox.

FE-1 Nikolai Budarin, in the Russian segment (RS), had a "full plate" 
today.  First, in support of the on-going upgrading of the Russian 
onboard computer system (BVS) to the new Vers. 7.01 software, he 
enabled a command response test by MCC-M of the integrated control 
panel (InPU) via S-band.

Budarin then was scheduled to install two new cables of the SUBA 
onboard equipment control system in the SM, brought up by 
10P/Progress-247, to reconfigure control of both SKV-1 and SKV-2 air 
conditioners. [During the actual connector matings and dematings on 
the systems, on MCC-M go-ahead, both SKVs were powered down.]

Budarin also reconfigured the Russian Ethernet onboard comm/data 
network, which is linked with the USOS OpsLAN (US segment/operations 
local area network) through the FGB via SSC (station support computer) 
Router #1 laptop and coaxial cable. [Budarin's replaced coax cables 
between RS laptops (via the BRI SmartSwitch router) with more advanced 
"twisted pair" cables and removed some older Ethernet network cables 
of its stage 1 config.]

Subsequently, after Bowersox had upgraded the SSC Router 1 laptop to 
the new software, he and Nikolai tested the reconfigured RS/USOS 
integrated network,. [Their checkout featured a "Ping" interaction 
test to check the connection between BRI-connected laptops and their 
links with USOS laptops, and a print test to check printing from 
different types of laptops to both printers.]

From the FGB, Budarin removed another no-longer-required electronics 
block (M34-17, A9) of the SUD motion control system, then took digital 
photographs of the exposed area for later downlink via OCA. [The 
images are required to help establish the configuration of containers 
under development for installation in these areas.]

Nikolai also installed three new SUBA cables on the SM command 
processor system (BOK) to provide capability for automatic 
(autonomous) control of the US matching unit for the Progress vehicle.

In the DC-1 "Pirs" docking module, Budarin performed the 
long-term/periodic task of verifying the status of circuit breakers on 
the BVP panel and fuses in the fuse box panel (BPP). Then, in the FGB 
module, he completed the monthly preventive task of replacing the 
filter cartridges in the PS1 and PS2 dust collectors of Zarya's 
ventilation system.

Ken Bowersox set up and activated the IV-CPDS (intravehicular charged
particle directional spectrometer), for taking EVARM (EVA radiation
monitoring) background radiation readings in the cabin. After 
subsequent data transfer to the computer, the equipment was stowed 
again.

Bowersox performed the scheduled monthly maintenance on the CSA-CP 
(compound specific analyzer-combustion products) which involves one 
hour of unattended air sample data collection time.

FE-2/SO Don Pettit continued his assistance in the ground's MSG
(microgravity science glovebox) troubleshooting, which further 
explored the electrical circuit anomaly by adding an additional power 
path. [Pettit again rotated the rack in order to make the necessary 
connection on the MSG electronics box. After Don's activations, the 
ground conducted extensive commanding and corresponding Ku-band 
telemetry checks. Purpose of these steps is to systematically track 
down and pinpoint the root cause behind the recent repeated PDC (power 
distribution controller) 12V and 5Vdc circuit breaker trip.]

Pettit completed the daily routine maintenance of the SM's SOSh life 
support system (including ASU toilet system), while the CDR prepared 
the regular IMS inventory delta file for downlink and performed the 
daily Lab payload status checkup on the PCG-STES010 payload.

All crewmembers performed their normal daily physical exercise on TVIS
(treadmill with vibration isolation and stabilization), RED (resistive
exercise device), CEVIS (cycle ergometer with vibration isolation) and 
VELO cycle with load trainer. [Since CEVIS has currently a problem 
with manual pre-selection of the loads required for the accelerated 
EVA prebreathing protocol, preferred prebreathing method is either 
in-suit or via Airlock "camp-out". However, CEVIS is still usable for 
general exercise.]

Bowersox and Pettit performed another session with the psychological 
MedOps WinSCAT (Spaceflight Cognitive Assessment Tool) experiment. 
[This is a time-constrained questionnaire test of cognitive abilities, 
routinely performed by astronauts aboard the ISS every 30 days before 
or after the PHS (periodic health status) test or on special CDR's, 
crewmember's or flight surgeon's request.]

Bowersox and Pettit spent time at the RWS (robotics workstation) 
reviewing the DOUG (dynamic operations ubiquitous graphics) software 
setup for tomorrow's operations with SSRMS/Canadarm2. The arm is 
currently based on PDGF1 (power and data grapple fixture #1) on the 
MBS (mobile base system), with its Tip LEE (latching end effector) 
grappled to PDGF2. [SSRMS operations tomorrow and on 3/6 will be used 
to perform surveys of the S1 and P1 truss segments. On S1, seven RBVM 
QD (radiator beam valve module quick disconnect) thermal covers and 
the zenith S0-S1 interface will be inspected.  On P1, two RBVM QD 
thermal covers and the port SVS (space vision system) target array 
will be inspected. After the S1 survey tomorrow, SSRMS will remain 
grappled to the Lab PDGF. If there is sufficient time, the MBS PDGF 1 
may be released and the arm maneuvered to the P1 RBVM survey position; 
if time does not allow it, the SSRMS reconfig for the P1 survey will 
occur on 3/6. Analysis (in the absence of telemetry data) has 
indicated that at least one thermal cover/bootie on the P1 RBVM 
appears to be not correctly positioned. If booties are not properly in 
place, the resulting higher temperatures over longer periods could 
potentially degrade the ammonia QD seals underneath.]

On POC (Payload Operations Center) request, Don Pettit powered up the 
HRF (Human Research Facility) and activated the GASMAP (gas analyzer 
system for metabolic analysis physiology) in standby mode. [This gives 
the GASMAP ion pump time to scrub out extra helium accumulated during 
the recent PuFF sessions.]

Yesterday the onboard C&W (caution & warning) system reported the loss 
of subset 1 of the Service Module (SM) TsVM Central Computer system, 
reducing redundancy in the Russian systems, with the other two lanes 
functioning OK.  [All three U.S. C&C MDMs (command & control 
computers) are nominal, and since the primary C&C is in sync with the 
primary subset #2 of TsVM, no action was required from ground or crew 
for the USOS C&DH (command & data handling) system at this time. The 
SM's TVM Terminal Computer has also been operating on two of three 
subsets since last year. Both TsVM and TVM are to be reinitialized 
with the new software on all three subsets this week, TVM tonight 
during crew sleep, TsVM tomorrow night.]

The software patch to correct the RGA (rate gyro assembly) "data 
staleness" problem was uplinked to both GNC (guidance, navigation & 
control) MDMs yesterday. While running overnight, unexpected control 
torques were observed in the ISS attitude control subsystem (ACS) due 
to stale RGA data. This anomaly was not observed prior to the patch 
installation. The motion control system (MCS) is now again using 
Russian attitude, angular rate, and state vector (position, velocity) 
data to preclude any anomalous RGA data propagating through the 
system.

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.